Startup NIGHT 11 - 20170912 UT

Tucson: Jenny, Olga, John Hill, Amali

Off site: Greg

Mountain: Pat

Summary: The Poptart worked. Seeing stable and nice, so we were able to obtain AGw transform data. Tests with LUCI1 AO were unsuccessful. Greg investigating. Amali worked on DX Adsec communications and some SX adsec tasks. We wrapped up with Multi-Slit and MOS tellurics, finishing with unsuccessful attempts at a MOS science OB. The acquisition of the first failed, when we didn\x92t see alignment stars we moved to a target with bright unambiguous alignment stars. The alignment boxes failed to appear when running the RTD MOS align routine. We attempted some basic troubleshooting before closing for the night (see IT 6777). Closed up and handed over to Dave for flexure verification.

At ~10:50 we lost network to outside connections for about 5 min, but maintained connection to the mountain. Lost to abell and google (log)

Night priorities outlined: https://docs.google.com/spreadsheets/d/1q3hQ04UESRPDJTzQn5SQ1EOXBOee_conayY7WJYgpRQ/edit#gid=0

18:40 Opened and sending test preset. DX is not cooperating.

18:58 Preset successful, John is working on collimating

19:07 light cirrus overhead, might put a stop to flats before they start.

19:11 IIF lies, working on getting the adsecs happy.

Some Adsec issues, Doug and John resolved

19:54 Seeing looks good! 0.7 on SX guider. Doug is composing a suitable list and we are setting up on stone L1 target for AGW transform.

20:10 First line of Doug\x92s script failed, broke co-pointing limit. Talked to him and it isn\x92t doing what he expected so he is checking it out in the simulator. He noticed some wavelength discrepancies too.

20:12 Seeing 0.7 on the guiders

Taking a couple darks while we wait.

Luci1.20170912.0026-30 0s

Luci2.20170912.0001-05 0s

Luci1.20170912.0031-35 6s

Luci2.20170912.0006-10 6s

20:31 Kinks worked out, starting L1, seeing 0.68\x94 on guider

Clear, H N30 Field Stop, Flexure on,

Luci1.20170912.0036

Luci2.20170912.0011

20:45 another geirs sync error. Solution in IT did not work. Claims not in eng mode so cannot perform put (did stop and restart geirs to ensure that set as master and still could not). KMALLOC_Split came up with the get as 8.0

20:54

Luci1.20170912.0053 0s

Luci2.20170912.0028 0s

17 luci1 0055 luci2 0030 Seeing 0.8\x94 on guider

60 luci1 0056 luci2 0031

124 luci1 0057 luci2 0032

156 luci1 0058 luci2 0033

166 MIssed guide star. Picked another one on right. Stopped to recenter pointing.

Start=5 (166) luci1 0059 luci2 0034

146 luci1 0060 luci2 0035

132 luci1 0061 luci2 0036

112 luci1 0062 luci2 0037 Seeing up to 1.12 on guider

167 luci1 0063 luci2 0038

101

52

Data no good. LUCI images do not look good and need upgrade

FeII images look elongated for luci1. Focus offset issue? Going back to H. Not collimated. John is resending.

21:46 Restarting StoneL1 transform with FeII 5x5s Seeing 0.83\x94 on guider. Mag 12-14, getting 20-25K counts

Collimating

17 luci1 0072 luci2 0044 (images smeared???)

60 luci1 0073 luci2 0045 images ok

124 luci1 0074 luci2 0046 collimation is poor again on LUCI1, LUCI2 ok

Collimated

156 luci1 0075 luci2 0047

166 luci1 0076 luci2 0047 image names?? Somethign odd

146 luci1 0077 luci2 0048

Collimated

132 luci1 0078 luci2 0049

112 luci1 0079 luci2 0050

167 luci1 0080 luci2 0051

Coll

101 luci1 0081 luci2 0052, LUCI1 elongated, 0.8\x94 on guider

52 luci1 0082 luci2 0053

44 luci1 0083 luci2 0054

Coll (note luci2 is taking more iterations to collimate than luci1 so luci1 is almost a full set ahead?)

16 luci1 0084 luci2 0055

13 Luci1 0085 ugly luci2 0055

87 luci1 0086 luci2 0057

Coll

160 luci1 0087 luci2 0058

105 luci1 0088 luci2 0059

159 luci1 0089 luci2 0060 on image 89, FWHM~5pix (0.625\x94) and on image 60 ~6pix( 0.85\x94)

Coll

126 luci1 0090 luci2 0061

42 luci1 0091 luci2 0062

110 luci1 0092 luci2 0063 LUCI2 a bit elongated

Coll

32 luci1 0093 luci2 0064 LUCI1 very poor image quality

55 luci1 0094 luci2 0065 LUCI1 looks round and 5.3 pix (0.67\x94)

68 luci1 0095 luci2 0066 0.88\x94 on guider

Coll

77 luci1 0096 luci2 0067

@22:48-22:51 lost internet connection to external sites (gmail, abell, etc, )

162 luci1 0097 luci2 0068

173 luci1 0098 luci2 0069

Coll

43 luci1 0099 luci2 0070

54 luci1 0100 luci2 0071

131 luci1 0101 luci2 0072

Coll

22 luci1 0102 luci2 0073

72 luci1 0103 luci2 0074 Seeing looks even better

139 luci1 0104 luci2 0075

Coll

7 luci1 0105 luci2 0076

93 luci1 0106 luci2 0077

At end for LUCI2 the image did not take automatically so in track for a long time for 0078

LUCI2 about 10 min behind LUC1

23:12 Seeing has been stable and good. John recommends a look at another field since this one is so low and we have had marginal images at times. We are going to go to Stone_P1. Contacting Greg for SX AO work

23:17 obs4 session insanely slow. Moving to obs2

Greg was not aware he was doing the AO night time checkout. Olga is making scripts to check on and off axis AO stars with different binnings. WHile prepping continuing with plan of collecting stone P data

23:27 Starting Stone_P1_50

Track Sky Luci1 0108 luci2 0080

Luci1 0109 luci2 0081

11 Luci1 0110 luci2 0082

10 Luci1 0111 luci2 0083

4 GS is a binary Luci1 0112 luci2 0084

Coll

2 Luci1 0113 luci2 0085

12 Luci1 0114 luci2 0086

9 Luci1 0110 luci2 0087

Coll

1 Luci1 0116 luci2 0088

7 Luci1 0117 luci2 0089

16 Luci1 0118 luci2 0090 Seeing 0.7\x94 on guider

Coll

6 Luci1 0119 luci2 0091

Coll

Track Luci1 0120 luci2 0092

23:56 Starting Stone_P1_65

Track Sky Luci1 0121 luci2 0093

Luci1 0122 luci2 0094

10 Luci1 0123 luci2 0095

4 Luci1 0124 luci2 0096

9 Luci1 0125 luci2 0097

Coll

7 Luci1 0126 luci2 0098

16 Luci1 0127 luci2 0099

6 Luci1 0128 luci2 0100

Coll

3 Luci1 0129 luci2 0101

8 Luci1 0130? luci2 0102 LUCI1 script stalled on GS8 so just retook GS8 in script 131

coll

Track Luci1 0132 luci2 0103

LUCI1 and LUCI2 collimated and ran script at the same rate for the P1 fields

LOGS were copied and are in /home/data/backuplogs as well as original location

0:28 Greg is getting things set up

0:43 Sending LUCI1 AO script, 12.1 mag AO ref star so should be bin 3

0:59 Secondary and WFS are not talking. John turned off LBTI WFS, and greg power cycled..

Cirrus wafting through

1:35 Greg is fighting with SX, Amali is working starting work with DX

1:44 Greg is giving up for now, Amali gets free reign with both adsecs.

2:15 Thin cloud moving through

2:43 Amali finished her tests. Reconfiguring for LUCI1+LUCI2.

John tried to implement new pointing models.

3:04 LUCI1+LUCI2 long-slit acq on telluric HIP 9116. G200HK, cwl 1.93, 1.5\x94 slit
LUCI 133-134 LUCI2 104-105

3:17 Acq done LUCI1 135 LUCI2 106 confirmation image Seeing 0.75 on guider

HKspec, cwl 1.93 Tell LUCI1 136-139 LUCI2 106-110

3:36 MOS Tell with M31NUC mask on HIP 10316 (moving around a little for pointing model). zJspec1.17 Acq Images LUCI1 140-142 LUCI2 111-113, Tell LUCI1 143-148 LUCI2 114-119

4:03 Slewing to M31 for MOS acq and science Bino OB. Think this is an ARGOS mask... Decided to try another target

4:30 p13Mask1, NO Reference star boxes when running MOS align. Not sure whats happening. Tied stopping and restarting aladdin, reloading mask. The alignment stars are near their boxes, so the mask is aligned not rotated or too out of whack, but not having the visible boxes show in teh RTD routine prevented us from doing the MOS alignment. Wrote up IT 6777

5:10 Finished troubleshooting. closed up and handed over to Dave for LUCI flexure verification

Comments

 
Topic revision: r2 - 12 Oct 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback